### v1

@@ A.1.2 Logical Reasoning Process
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\frac{\begin{array}[]{c}loc\_in(s_{1},o_{1}),\leavevmode\nobreak\ loc\_in(s_{2},o_{2}),\leavevmode\nobreak\ o_{1}=s_{2}\\ \end{array}}{\begin{array}[]{c}loc\_in(s_{1},o_{2})\end{array}}[Trans].$
@@
Composite Reasoning Rules $[Comp]$. Given multiple basic reasoning rules or predicates $[Rule_{i}]\in\{[Neg],[Sym],[Inverse],[Trans],[Predicates]\}$, we can chain them up to form a new composite reasoning rule.
[…]
@@ A.3.2 Facts Similarity Algorithm
  • Given evidence triples $[(A,\_,B),(A,\_,C)]$ and response triples $[(A,\_,B^{\prime}),(A,\_,C^{\prime}),(A,\_,D)]$
@@
  • Store all similarity scores in $similarity\_list$
@@
  • Given evidence triples $[(A,\_,B),(C,\_,D)]$ and response triples $[(A,\_,B^{\prime}),(C,\_,E)]$
@@ B.4 Summary Statistics
  • Notes: SDs are in parentheses. All biases are significantly different from zero ($p<0.0001$)
@@ C.2 Summary Statistics
Pre-test Post-test Pre-test Post-test $N$
@@ C.3 Subsample Analysis of Treatment Effects
$\Delta$Accuracy $\Delta$Bias
@@
  • Notes:${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$. Standard errors in parentheses are clustered at the player and question levels. The reference category is the group with high LLM confidence and high human baseline confidence.
@@ C.4 Gender Gap in Overconfidence
Appendix D The Results in the Main Text
@@ D.1.1 Description of Table 1
$N$ 9,779 9,895 9,974 9,614 9,854
@@ D.2.2 Comparison of Confidence Measures
$N$ 9,779 9,895 9,974 9,614 9,854
R2 0.003 0.06 0.07 0.05 0.02
  • Notes: Dependent variable is correct answer (= 1). Confidence levels are normalized by subtracting 1 from the original confidence scores, so the constant term represents accuracy at 100% confidence ($\tilde{p}=1$). Regression is weighted by the frequency of confidence levels. All coefficients are significant at p<0.01.
@@ D.3.2 Confidence Gradients on Human Benchmark Subset
Confidence levels are normalized by subtracting 1 from the original confidence scores, so the constant term represents accuracy at 100% confidence ($\tilde{p}=1$). These regression estimates are weighted by the frequency of observations at each confidence level. For human participants, standard errors are clustered at both individual and question levels. All reported coefficients are statistically significant at p < 0.01.
Table S11: The confidence gradient in accuracy (Human Baseline Subsample)
human gpt_3.5 gpt_4o gpt_o1 llama3.1_8b llama3.2_3b
Confidence 0.52 1.49 2.65 2.61 1.55 0.42
@@
$N$ 5,220 1,933 1,994 1,997 1,958 1,984
$R^{2}$ 0.05 0.02 0.09 0.08 0.11 0.003
@@ D.4.1 Description of Figure 2
The key variables of interest in the LLM-exposure experiment are the differences in accuracy and bias (measured by the gap between self-reported confidence and actual correctness) between Pre-treatment and post-treatment measurements.Table S12 presents regression results across treatment conditions, where $\Delta$accuracy and $\Delta$bias represent the differences between baseline and intervention stages for each respective measure.
@@
$\Delta$Accuracy $\Delta$Bias
@@ D.5 Table 3: Subsample Analysis of LLM Exposure Experiment
In table 3, we examine heterogeneity across two dimensions: (1) how the presence of LLM confidence information influences participants’ judgment and decision-making, (2) how participants’ initial confidence levels moderate the effects of LLM assistance. This analysis provides insights into the contexts and populations for whom LLM assistance offers the greatest benefit.
@@ D.7 Similarity Measure Results
D.8 Replication Results
D.8.1 Descriptive Summary
Table S17: Replication Results: Accuracy and Confidence Measures
a) Fraction of b) LLM confidence in … c) Derived
GPT 3.5
     1 0.35 0.94 0.96 0.91 0.87
(0.476) (0.045) (0.044) (0.061) (0.082)
     2 0.35 0.94 0.95 0.91 0.87
(0.476) (0.045) (0.044) (0.061) (0.081)
     3 0.35 0.94 0.95 0.91 0.87
(0.477) (0.045) (0.045) (0.061) (0.082)
     4 0.35 0.94 0.96 0.91 0.87
(0.477) (0.045) (0.043) (0.061) (0.082)
     5 0.35 0.94 0.96 0.91 0.87
(0.477) (0.045) (0.044) (0.061) (0.082)
     6 0.35 0.94 0.95 0.91 0.87
(0.476) (0.045) (0.045) (0.061) (0.082)
GPT 4o
     1 0.63 0.94 0.95 0.93 0.88
(0.482) (0.053) (0.044) (0.054) (0.083)
     2 0.65 0.93 0.94 0.92 0.87
(0.478) (0.058) (0.048) (0.055) (0.085)
     3 0.65 0.93 0.94 0.92 0.87
(0.478) (0.058) (0.049) (0.055) (0.085)
     4 0.64 0.93 0.94 0.92 0.87
(0.479) (0.059) (0.048) (0.054) (0.085)
     5 0.64 0.93 0.94 0.92 0.87
(0.479) (0.059) (0.049) (0.055) (0.086)
     6 0.65 0.93 0.94 0.93 0.87
(0.476) (0.055) (0.046) (0.052) (0.082)
Panel A: GPT-3.5
Panel B: GPT-4o
Figure S9: Proportion of Correctness Combinations Across Six Rounds
Table S18: Differences in Accuracy and Confidence: GPT-3.5
Accuracy Confidence
Rounds Mean (SD) p-value Mean (SD) p-value
1_2 0.001 (0.238) 0.733 -0.000 (0.040) 0.810
1_3 0.002 (0.236) 0.439 -0.000 (0.039) 0.469
1_4 0.002 (0.239) 0.309 -0.000 (0.040) 0.291
1_5 0.002 (0.238) 0.395 -0.000 (0.039) 0.378
1_6 -0.002 (0.258) 0.346 0.000 (0.042) 0.998
2_3 0.001 (0.219) 0.611 -0.000 (0.039) 0.622
2_4 0.002 (0.216) 0.372 -0.000 (0.039) 0.397
2_5 0.001 (0.226) 0.822 -0.000 (0.040) 0.531
2_6 -0.003 (0.254) 0.263 0.000 (0.042) 0.809
3_4 0.001 (0.215) 0.572 -0.000 (0.039) 0.725
3_5 -0.001 (0.224) 0.751 -0.000 (0.040) 0.885
3_6 -0.005 (0.257) 0.076 0.000 (0.042) 0.490
4_5 -0.001 (0.217) 0.708 0.000 (0.039) 0.836
4_6 -0.005 (0.258) 0.065 0.000 (0.042) 0.311
5_6 -0.004 (0.257) 0.114 0.000 (0.042) 0.410
Table S19: Differences in Accuracy and Confidence: GPT-4o
Accuracy Confidence
Rounds Mean (SD) p-value Mean (SD) p-value
1_2 0.012 (0.213) $0$0.006 (0.037) 0.000
1_3 0.012 (0.216) $0$0.006 (0.037) 0.000
1_4 0.010 (0.215) $0$0.007 (0.036) 0.000
1_5 0.009 (0.211) $0$0.006 (0.039) 0.000
1_6 0.017 (0.238) $0$0.003 (0.038) 0.000
2_3 -0.000 (0.096) 0.917 -0.000 (0.029) 0.760
2_4 -0.002 (0.116) 0.143 -0.000 (0.031) 0.183
2_5 -0.002 (0.112) 0.033 -0.000 (0.030) 0.796
2_6 0.005 (0.189) 0.007 0.003 (0.035) 0.000
3_4 -0.002 (0.117) 0.170 -0.000 (0.028) 0.259
3_5 -0.002 (0.107) 0.032 0.000 (0.029) 0.975
3_6 0.005 (0.195) 0.008 0.003 (0.036) 0.000
4_5 -0.001 (0.114) 0.538 0.000 (0.030) 0.273
4_6 0.007 (0.199) 0.001 0.003 (0.037) 0.000
5_6 0.008 (0.196) 0.000 0.003 (0.037) 0.000
D.8.2 Regression Analysis of Response Persistence
Model 1 (columns 1-2) examines the persistence in response accuracy across rounds. The strong coefficients on previous accuracy (0.8830 for GPT-3.5 and 0.9458 for GPT-4o, both significant at the 0.1% level) indicate high consistency in whether responses are correct across rounds.
Model 2 (columns 3-4) analyzes how previous accuracy influences confidence levels in subsequent rounds. While statistically significant, the small coefficients (0.0113 for GPT-3.5 and 0.0305 for GPT-4o) suggest that previous accuracy has limited impact on future confidence levels. The high constant terms (0.9384 and 0.9384) indicate that both models maintain high baseline confidence levels regardless of previous performance.
Model 3 (columns 5-6) combines both effects by including both previous confidence and previous accuracy on current confidence. The results reveal strong persistence in confidence levels, particularly for GPT-4o (coefficient of 0.8365 compared to 0.6035 for GPT-3.5). Once controlling for previous confidence, the effect of previous accuracy becomes small, suggesting again that confidence patterns are largely independent of actual performance.
Table S20: Response Persistence and Round Effects
Response Correctness Confidence Level Combined Analysis
GPT-3.5 GPT-4o GPT-3.5 GPT-4o GPT-3.5 GPT-4o
Previous Accuracy 0.8830*** 0.9458*** 0.0113*** 0.0305*** 0.0041*** 0.0046***
(0.0035) (0.0021) (0.0007) (0.0012) (0.0003) (0.0003)
Previous Confidence 0.6035*** 0.8365***
(0.0092) (0.0069)
Round 3 0.0004 -0.0112*** -0.0002 -0.0004* -0.0002 0.0051***
(0.0037) (0.0024) (0.0004) (0.0003) (0.0006) (0.0005)
Round 4 0.0006 -0.0127*** -0.0004* -0.0008* -0.0002 0.0048***
(0.0030) (0.0024) (0.0004) (0.0003) (0.0005) (0.0004)
Round 5 -0.0014 -0.0119*** -0.0003* -0.0004* -0.0000 0.0055***
(0.0031) (0.0023) (0.0004) (0.0003) (0.0005) (0.0004)
Round 6 -0.0049* -0.0036* -0.0000 0.0026*** 0.0002 0.0081***
(0.0034) (0.0030) (0.0004) (0.0004) (0.0005) (0.0005)
Question Order $0**$$0$$0$$0**$$0$0.0000*
(0.0000) (0.0000) (0.0000) (0.0000) (0.0000) (0.0000)
Constant 0.0360*** 0.0463*** 0.9384*** 0.9136*** 0.3720*** 0.1446***
(0.0029) (0.0029) (0.0006) (0.0012) (0.0087) (0.0064)
Num.Obs. 48588 49996 48986 49994 48983 49965
R2 0.781 0.898 0.015 0.065 0.372 0.711
  • Notes: ***Significant at the 1 percent level; **Significant at the 5 percent level; *Significant at the 10 percent level.

### v2

@@ A.1.2 Logical Reasoning Process
$\frac{\begin{array}[]{c}\mathit{nm}(s,o)\\ \end{array}}{\begin{array}[]{c}\neg\penalty 10000\ \overline{\mathit{nm}}(s,o)\end{array}}[Neg]$
An example of this type of rule is: $\frac{\begin{array}[]{c}\mathit{was}(s,o)\\ \end{array}}{\begin{array}[]{c}\mathit{\neg\penalty 10000\ wasn^{\prime}t}(s,o)\end{array}}[Neg]$.
@@
$\frac{\begin{array}[]{c}\mathit{nm}(s,o)\\ \end{array}}{\begin{array}[]{c}\mathit{nm}(o,s)\end{array}}[Sym]$
An example of this type of rule is: $\frac{\begin{array}[]{c}\mathit{different\_from}(s,o)\\ \end{array}}{\begin{array}[]{c}\mathit{different\_from}(o,s)\end{array}}[Sym]$.
@@
$\frac{\begin{array}[]{c}\mathit{nm}(s,o),\mathit{nm}^{\prime}=Reverse(\mathit{nm})\\ \end{array}}{\begin{array}[]{c}\mathit{nm}^{\prime}(o,s)\end{array}}[Inverse]$
An example of this type of rule is: $\frac{\begin{array}[]{c}\mathit{influence\_by}(s,o)\end{array}}{\begin{array}[]{c}\mathit{influence}(o,s)\end{array}}[Inverse]$. With this rule, from the triple (Haruki Murakami, influence_by, Richard Brautigan), we can derive a new triple (Richard Brautigan, influence, Haruki Murakami).
@@
$\frac{\begin{array}[]{c}\mathit{nm}(s_{1},o_{1}),\penalty 10000\ \mathit{nm}(s_{2},o_{2}),\penalty 10000\ o_{1}=s_{2}\\ \end{array}}{\begin{array}[]{c}\mathit{nm}(s_{1},o_{2})\end{array}}[Trans]$
@@
$\frac{\begin{array}[]{c}loc\_in(s_{1},o_{1}),\penalty 10000\ loc\_in(s_{2},o_{2}),\penalty 10000\ o_{1}=s_{2}\\ \end{array}}{\begin{array}[]{c}loc\_in(s_{1},o_{2})\end{array}}[Trans].$
@@
Composite Reasoning Rules $[Comp]$. Given multiple basic reasoning rules or predicates $Rule_{i}]\in\{[Neg],[Sym],[Inverse],[Trans],[Predicates]\}$, we can chain them up to form a new composite reasoning rule.
$\frac{\frac{\frac{\begin{array}[]{c}\mathit{nm}_{1\_Rule_{1}}(...),\mathit{nm}_{2\_Rule_{1}}(...),...\\ \end{array}}{\begin{array}[]{c}R_{1}\end{array}}[Rule_{1}],\;...}{\frac{\frac{\begin{array}[]{c}...\end{array}}{\begin{array}[]{c}...\end{array}}[...],\;...}{\frac{\begin{array}[]{c}\mathit{nm}_{1\_Rule_{i}}(...),\mathit{nm}_{2\_Rule_{i}}(...),...\\ \end{array}}{\begin{array}[]{c}R_{\mathit{i}}\end{array}}[Rule_{i}],\;...}}}{\begin{array}[]{c}R_{\mathit{new}}\end{array}}[Comp]$
@@ A.3.2 Facts Similarity Algorithm
  • Given evidence triples [(A,_,B),(A,_,C)][(A,\_,B),(A,\_,C)] and response triples $(A,\_,B^{\prime}),(A,\_,C^{\prime}),(A,\_,D)]$
@@
  • Store all similarity scores in similarity_listsimilarity\_list
@@
  • Given evidence triples [(A,_,B),(C,_,D)][(A,\_,B),(C,\_,D)] and response triples $(A,\_,B^{\prime}),(C,\_,E)]$
@@ B.4 Summary Statistics
  • Notes: SDs are in parentheses. All biases are significantly different from zero ($p<0.001$)
@@ C.2 Summary Statistics
Pre-test Post-test Pre-test Post-test NN
@@ C.3 Subsample Analysis of Treatment Effects
$\DeltaAccuracy Δ\DeltaBias$
@@
[…]
@@ C.4 Gender Gap in Overconfidence
Appendix D The Results from the Exhibits in the Main Text
@@ D.1.1 Description of Table 1
NN 9,779 9,895 9,974 9,614 9,854
  • Notes: All overconfidence measures are significantly different from zero at p<0.001. Confidence levels are normalized by subtracting 1 from the original confidence scores.
@@ D.2.2 Comparison of Confidence Measures
NN 9,779 9,895 9,974 9,614 9,854
R2 0.003 0.07 0.06 0.05 0.02
  • Notes: Dependent variable is correct answer (= 1). Confidence levels are normalized by subtracting 1 from the original confidence scores, so the constant term represents accuracy at 100% confidence ($\tilde{p}=1$). All coefficients are significant at p<0.001.
@@ D.3.2 Confidence Gradients on Human Benchmark Subset
Confidence levels are normalized by subtracting 1 from the original confidence scores, so the constant term represents accuracy at 100% confidence ($\tilde{p}=1$). These regression estimates are weighted by the frequency of observations at each confidence level. For human participants, standard errors are clustered at both individual and question levels.
Table S11: Confidence Gradient in Accuracy (Human Baseline Subsample)
human gpt 3.5 gpt 4o gpt o1 llama3.1 llama3.2
Confidence 0.51 1.39 2.50 2.26 1.42 0.32
@@
NN 5,220 1,933 1,994 1,997 1,958 1,984
$R^{2}$ 0.05 0.02 0.09 0.08 0.11 0.002
  • Notes: Dependent variable is correct answer (= 1). The table presents the gradient of accuracy with respect to self-reported confidence in answer correctness across models using the human baseline subsample. Confidence levels are normalized by subtracting 1 from the original confidence scores. All coefficients are significant at p<0.001, except for Llama 3.2, which is significant at p<0.05.
@@ D.4.1 Description of Figure 2
[…]
@@
$\DeltaAccuracy Δ\DeltaBias$
@@ D.5 Table 3: Subsample Analysis of LLM Exposure Experiment
In table 3, we examine heterogeneity across two dimensions: (1) how the presence of LLM confidence information influences participants’ judgment and decision-making, (2) how participants’ initial confidence levels moderate the effects of LLM assistance. The full regression results see Table S13 below. This analysis provides insights into the contexts and populations for whom LLM assistance offers the greatest benefit.
Table S13: Subsample Analysis of LLM Exposure Experiment - Full Results
Human Baseline Confidence LLM Confidence
Dependent variable: $\Delta$ Accuracy $\Delta$ Bias $\Delta$ Accuracy $\Delta$ Bias
LLM Answer 0.086 0.070 -0.033 0.133
(0.015) (0.017) (0.018) (0.020)
LLM Answer + Conf. 0.119 0.141 -0.028 0.160
(0.017) (0.018) (0.020) (0.021)
High Confidence (Human) 0.001 -0.018
(0.005) (0.005)
LLM Answer $\times$ High Confidence (Human) -0.064 -0.063
(0.018) (0.019)
LLM Answer + Conf. $\times$ High Confidence (Human) -0.091 -0.124
(0.019) (0.019)
High Confidence (LLM) 0.008 -0.008
(0.005) (0.006)
LLM Answer $\times$ High Confidence (LLM) 0.136 -0.138
(0.020) (0.021)
LLM Answer + Conf. $\times$ High Confidence (LLM) 0.153 -0.119
(0.024) (0.024)
Constant 0.003 0.018 -0.002 0.014
(0.004) (0.005) (0.004) (0.005)
Observations 11,610 11,610 10,957 10,957
R2 0.01082 0.01700 0.02477 0.01915
  • Notes: The table presents the full regression results including all control variables and interaction terms. The main text Table 3 presents a condensed version focusing on key treatment interactions.
  • $^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$.
@@ D.7 Similarity Measure Results
  • Notes: ∗∗∗p << 0.001; ∗∗p << 0.01; p << 0.05.